# Representative Action Selection for Large Action Space Meta-Bandits

### v3

@@ Abstract
We study the problem of selecting a subset from a large action space shared by a family of bandits, with the goal of achieving performance nearly matching that of using the full action space. We assume that similar actions tend to have related payoffs, modeled by a Gaussian process. To exploit this structure, we propose a simple $\epsilon$-net algorithm to select a representative subset. We provide theoretical guarantees for its performance and compare it empirically to Thompson Sampling and Upper Confidence Bound.
@@ 1 Introduction
We study a family of bandits that share a common but extremely large action space. We aim to understand whether it is possible—and how—to select a smaller set of representative actions that performs nearly as well as the full action space across all bandit instances. To build intuition, imagine a pharmacy preparing its inventory for the upcoming season. The available drugs (actions) are nearly infinite, and each customer (bandit) has unique characteristics. If two drugs share similar ingredients, their effects on a patient are likely to be similar. Likewise, if two patients have comparable health indices, a drug is likely to have similar effects on both. By modeling the expected outcome of each drug for each patient as a Gaussian process, we can capture these correlations. This correlation structure makes subset selection particularly interesting: if two drugs treat the same disease, choosing one may suffice, while drugs for distinct diseases might both be kept. In addition, demand also matters: drugs for rare illnesses can be excluded to save space, while flu medications should be stocked during flu season.
Different from prior approaches in multi-armed bandits (MAB) (lai1985asymptotically, ) that aim for identifying either a single best action or a subset that achieves high cumulative outcomes for a fixed bandit, our objective focuses on selecting a subset that is likely to contain the best action, or one whose best element performs nearly as well for a family of bandits. This problem can be seen as a large-scale combinatorial optimization under uncertainty, with applications where decisions involve a vast number of possibilities but are constrained by computational or time limitations for evaluating all options, e.g., inventory management, online recommendations.
Consider the following setting: In a bandit, if a decision-maker plays an action with a fixed but unknown feature vector $a\in\mathcal{A}_{\operatorname*{full}}\subset\mathbb{R}^{n}$, they observe a random outcome taking values in $\mathbb{R}$. We define the expected outcome of playing action $a$ in this bandit as $\mu_{a}(\theta):=\langle a,\theta\rangle$ where this bandit instance $\theta\in\mathbb{R}^{n}$ is drawn from an unknown multivariate Gaussian distribution. Thus, the collection of random variables $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ forms a Gaussian process (GP) (vershynin2018high, , Chapter 7), while a more general sub-Gaussian assumption is considered in the Supplementary. Our setting aligns with that of contextual bandits (dani2008stochastic, ), and the restriction to linear functions may be partially mitigated by allowing $n$ to be unbounded.
@@
Consider that the decision-maker has access to a fixed action subset $\mathcal{A}\subset\mathcal{A}_{\operatorname*{full}}$. For a given bandit instance $\theta$, if the optimal action over the full space lies within $\mathcal{A}$, the decision-maker benefits from reduced suboptimal actions to explore. Conversely, if the optimal action lies outside $\mathcal{A}$, regret arises from being unable to select this best action. Thus, for a bandit instance $\theta$, we define the regret as the difference in expected outcome between having access to the full action space versus being restricted to the subset:
$\boxed{\textnormal{Regret}(\theta):=\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta)-\max_{a^{\prime}\in\mathcal{A}}\;\mu_{a^{\prime}}(\theta),}$ (1)
@@
which depends on the sampled bandit instance and is therefore a random variable. Our objective is to identify a small subset $\mathcal{A}$ that minimizes the expected regret $\mathbb{E}_{\theta}\left[\operatorname*{Regret}\right]$ over all possible bandit instances, making the underlying optimization both stochastic and combinatorial.
This objective is motivated by practical considerations. The classic Bayesian regret in the bandit literature (agrawal2012analysis, ) typically scales with the number of available actions. However, if a subset $\mathcal{A}$ is carefully chosen, the resulting Bayesian bandit regret can be significantly lower. This is especially beneficial when the action space is large, as even the initialization phase can be computationally expensive. To see this, we can decompose the bandit regret as follows:
@@
$\textnormal{BayesianBanditRegret}:=\mathbb{E}\sum_{t=1}^{N}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta)-\mu_{A_{t}}(\theta)\right]$
$=\mathbb{E}\sum_{t=1}^{N}\left[\max_{a\in\mathcal{A}}\mu_{a}(\theta)\!-\!\mu_{A_{t}}+\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta)\!-\!\max_{a\in\mathcal{A}}\mu_{a}(\theta)\right]$
$\leq C\sqrt{|\mathcal{A}|\cdot N\log N}+N\cdot\mathbb{E}_{\theta}[\textnormal{Regret}].$
In the first equality, $A_{t}\in\mathcal{A}$ denotes the action chosen by a policy, e.g., Thompson Sampling (agrawal2012analysis, ), in round $t$, and $N$ is the number of rounds. The expectation is taken over the randomness in the distribution of bandit instances and in actions selected by the policy. The inequality follows from the well-known regret bounds for Thompson Sampling (lattimore2020bandit, ) [^1] This bound could be relaxed if the feature vector $a$ is known, or if $a$ can be incorporated as the input of a kernel function, as discussed in related works. However, we use this general bound because we also consider the most general case where $\mathcal{A}_{\operatorname*{full}}$ is simply an index set, e.g., $\mathcal{A}_{\operatorname*{full}}:=\{\textnormal{treatment A},\textnormal{treatment B},\textnormal{treatment C}\}$., where $|\mathcal{A}|$ denotes the cardinality of the action set and $C>0$ is a constant. Note that if the policy has access to the full action space, the Bayesian bandit regret is instead bounded by $C\sqrt{|\mathcal{A}_{\operatorname*{full}}|\cdot N\log N}$. Our main contributions are organized as follows:
  • Meta-Bandits Framework. We propose a meta-bandits framework that specifically tackles combinatorial action selection by leveraging correlations across similar actions and bandit instances. To the best of our knowledge, this is the first such framework.
  • $\epsilon$-Net Algorithm. We introduce a simple algorithm within this framework. It starts with the intuitive idea of placing a grid over the action space, then refines it using an importance-based selection mechanism.
  • Regret Analysis. We provide theoretical guarantees for both the grid and the algorithm’s output, including upper and lower bounds on expected regret, along with results under general sub-Gaussian processes. We also discuss the cost of not using a grid, which depends on the importance-structure of the action space.
  • Generalization and Empirical Validation. We extend the analysis to settings where outcome functions are sampled from a reproducing kernel Hilbert space (RKHS), and empirically compare our algorithm to Thompson Sampling (TS) and Upper Confidence Bound (UCB).
1.1 Related Works
Multi-Armed Bandits (lai1985asymptotically, ; auer2002finite, ) is defined by a set of actions (arms), each deliver outcomes that are independently drawn from a fixed and unknown distribution. The decision-maker sequentially selects an action, observes its outcome, and aims to maximize cumulative outcomes over time. Popular methods include the UCB (auer2002finite, ), TS (agrawal2012analysis, ), and EXP3 (auer2002nonstochastic, ) for adversarial settings.
Optimal Action Identification focuses on identifying the action with the highest expected outcome in a MAB setting using as few samples as possible (jamieson2014best, ; kaufmann2016complexity, ). Popular methods in fixed confidence setting include Action Elimination even2006action ; karnin2013almost , UCB, and LUCB, all of which achieve sample complexity within a $\log(|\mathcal{A}_{\operatorname*{full}}|)$ factor of the optimum. In fixed budget setting, there is Successive halving karnin2013almost , successive reject audibert2010best .
Stochastic Linear Optimization assumes that the expected outcome of each action depends through the inner product between a context $\theta$ and an action $a\in\mathcal{A}_{\operatorname*{full}}$ auer2002using ; dani2008stochastic ; rusmevichientong2010linearly . This line of work assumes that the action feature vectors are known, so the cardinality of the action space does not play a role and may even be infinite. soare2014best study the sample complexity of optimal action identification in this setting, which scales linearly with the action space dimension $n$.
GP Optimization addresses the case where the feature vectors of actions are unknown but a kernel function is available, so the inputs to the kernel are required. srinivas2009gaussian model the outcome function as a sample from a GP prior with a kernel function (williams2006gaussian, ), achieving bandit regret that scales as $\sqrt{\log|\mathcal{A}_{\operatorname*{full}}|}$ for finite action space. The widespread adoption of this method in bandit settings (valko2013finite, ; li2022gaussian, ), as well as in continuous action spaces (chowdhury2017kernelized, ), highlights the practicality of assuming that action outcomes are correlated. This also motivates our extension, where the outcome function of each bandit instance is modeled as a sample from a RKHS.
@@ 1.1 Related Works
Before discussing two seemingly relevant lines of work that select a subset from the action space, we first highlight an advantage of our approach. These methods assume a fixed subset cardinality $|\mathcal{A}|=K$, which is often unclear in practice and requires restarting the algorithm when changed. In contrast, our algorithm can adapt the subset size on the fly.
@@
Top-K Action Identification aims to identify the $K$ actions with the highest expected outcomes using as few samples as possible (kalyanakrishnan2012pac, ; gabillon2012best, ; kaufmann2016complexity, ; chen2017nearly, ). This line of work assumes that all actions are independent and have distinct expected values, making its methods inapplicable to our framework. If one were to apply these methods regardless, the most reasonable approach, in our view, would be to treat the family of bandit instances as a super-bandit, where each bandit instance corresponds to a round, and the expected payoff of an action in that round is given by $\mu_{a}$. In this setting, top-$K$ identification would refer to selecting $K$ actions with the highest expected payoffs $\mathbb{E}[\mu_{a}]$. In contrast, our framework considers that the expected outcome of each action, averaged over the distribution of bandits, may be the same—i.e., $\mathbb{E}[\mu_{a}]\!=\!c$ for all $a\!\in\!\mathcal{A}_{\operatorname*{full}}$, where $c$ is a constant—so that the entire action space shares the same highest expected payoff. Further, even if $\mathbb{E}[\mu_{a}]$ varies across actions, ignoring correlations can be fatal in our framework:
Example 1.
@@
Combinatorial Bandits considers that the decision maker selects $K$ of base arms from $\mathcal{A}_{\operatorname*{full}}$ in each round, forming a super arm $\mathcal{A}$, with $|\mathcal{A}|=K$. Popular methods include CUCB chen2016combinatorial , CTS wang2018thompson . We argue that this line of work is not directly applicable to our framework, but we include it as a baseline in subsequent empirical evaluations: (1) It assumes that the expected outcome of a super arm depends only on the expected outcomes of its individual base arms, or imposes a stricter monotonicity condition. In our case, even though $\mathbb{E}[\mu_{a}]=c$ for all $a\in\mathcal{A}_{\operatorname*{full}}$, super arm expected outcomes $\mathbb{E}[\max_{a\in\mathcal{A}}\mu_{a}]$ can differ significantly due to correlations among actions. (2) It assumes independence across base arms, whereas we explicitly model correlations. Ignoring these correlations misses the core challenge—an issue illustrated in Example 1.
2 Subset Selection Framework
We consider the problem of selecting a small number of representative actions from a large action space $\mathcal{A}_{\operatorname*{full}}\subset\mathbb{R}^{n}$, where $n\in\mathbb{N}$. (This framework applies to the case $n=+\infty$, with the additional assumption $\sum_{i\geq 1}a^{2}_{i}\!<\!\infty$.) The expected outcome depends on both the chosen action $a\in\mathbb{R}^{n}$ and an observed context $g\in\mathbb{R}^{n}$, and includes a constant $c\in\mathbb{R}$:
$\mu_{a}:=\langle a,g\rangle+c,\quad\forall\;a\in\mathcal{A}_{\operatorname*{full}},\;g\sim\mathcal{N}(0,\Sigma),$ (2)
where $\Sigma$ is a positive semi-definite matrix. Let $\theta\in\mathbb{R}^{n}$ follows a multivariate normal distribution with zero mean and identity covariance matrix $I$. The distribution of $g$ is in fact equivalent to $\Sigma^{1/2}\theta$. Now, let $\sigma_{j}$ denote the $j$-row of the matrix $\Sigma^{1/2}$. We have $g=(\langle\sigma_{j},\theta\rangle)_{j\leq n}$ and
$\langle a,g\rangle=\sum_{j\leq n}a_{j}\langle\sigma_{j},\theta\rangle=\left\langle\sum_{j\leq n}a_{j}\sigma_{j},\theta\right\rangle=\langle\Sigma^{1/2}a,\theta\rangle.$
Therefore, the setting in equation 2 is equivalent to $\mu_{a}:=\langle a,\theta\rangle+c,\;\forall\;a\in\Sigma^{1/2}\mathcal{A}_{\operatorname*{full}},\;\theta\sim\mathcal{N}(0,I),$ where $\Sigma^{1/2}\mathcal{A}_{\operatorname*{full}}$ denotes the image of $\mathcal{A}_{\operatorname*{full}}$ under the linear transformation $\Sigma^{1/2}$. Since the constant $c$ does not affect the regret (as defined in equation 1), we can, without loss of generality, focus on this canonical Gaussian process (vershynin2018high, , Chapter 7) in the remainder:

### v4

# Representative Action Selection for Large Action Space Bandit Families

@@ Abstract
We study the problem of selecting a subset from a large action space shared by a family of bandits, with the goal of achieving performance nearly matching that of using the full action space. Indeed, in many natural situations, while the nominal set of actions may be large, there also exist significant correlations between the rewards of different actions. In this paper we propose an algorithm that can significantly reduce the action space when such correlations are present, without the need to a-priori know the correlation structure. We provide theoretical guarantees on the performance of the algorithm and demonstrate its practical effectiveness through empirical comparisons with Thompson Sampling and Upper Confidence Bound methods.
@@ 1 Introduction
The multi-armed bandit (MAB) problem (Lai & Robbins, 1985; Auer et al., 2002a) models sequential decision-making under uncertainty: at each round, an agent chooses an arm from a fixed set and receives a stochastic reward drawn from an unknown distribution. The objective is to maximize cumulative reward by balancing exploration and exploitation.
In modern applications, action spaces are often extremely large, yet many related decision problems share this same set of actions. We ask a fundamental question: Can we select a small, representative subset of actions that performs nearly as well as the full action space across all such related problems? If possible, this would dramatically reduce computational and statistical complexity with only controllable performance loss. In this work, we formalize this question and investigate when and how such a small, universally effective action subset can be constructed.
As an example, consider a pharmacy preparing its inventory for the upcoming season. The set of available drugs will correspond to actions, and can be extremely large. Each patient will be represented by a bandit instance, i.e., a unique mapping from drugs to health outcomes based on their physiology and history. However, when multiple drugs contain similar ingredients, their effects on patients are likely to be similar. Therefore, we can retain just one or a few representative drugs from such a group without meaningfully reducing coverage or effectiveness.
In the classical multi-armed bandit problem (MAB) (Lai & Robbins, 1985; Lattimore & Szepesvári, 2020), the objective is to maximize cumulative reward over a time horizon $T$ for a single, fixed bandit. This is achieved by exploiting a single optimal action after exploration over the action space. In contrast, our objective is to select an action subset that likely contains near-optimal actions jointly for a family of bandits. This problem is a large-scale combinatorial optimization under uncertainty (Van Hentenryck & Bent, 2006; Li & Yuan, 2013).It arises in settings with too many options to evaluate exhaustively— such as route planning (Feng et al., 2025) and inventory management (Namir et al., 2022)—due to computational or time constraints.
Consider a Bayesian multi armed bandits setting, (Slivkins et al., 2019; Lattimore & Szepesvári, 2020): Let $\mathcal{F}$ be a family of bandits which share an action space $\mathcal{A}_{\operatorname*{full}}$, and let $\mathcal{P}$ be a probability distribution over $\mathcal{F}$. For a bandit instance $\theta\in\mathcal{F}$, let $\mu_{a}=\mu_{a}(\theta)$ denote the expected reward of an action $a\in\mathcal{A}_{\operatorname*{full}}$ given $\theta$. Each $\mu_{a}$ is then a random variable as $\theta$ is sampled from $\mathcal{P}$, and the collection $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ constitutes a stochastic process indexed by the action space.
In this paper we consider the similarity structure on the actions $\mathcal{A}_{\operatorname*{full}}$ that is induced by the correlations of the variables $\mu_{a}$. In this setting, the actions $a$ and $a^{\prime}$ will be considered similar if the rewards $\mu_{a}$ and $\mu_{a}^{\prime}$ are similar for most bandits sampled from $\mathcal{P}$. Note that this notion does not depend on the nominal properties of the set $\mathcal{A}_{\operatorname*{full}}$, such as the dimension in which $\mathcal{A}_{\operatorname*{full}}$ happens to be represented, but depends only on the statistical properties of the family of bandits $\mathcal{F}$ and the distribution $\mathcal{P}$. This provides a robust practical modeling framework which in turn may be used as the basis for the action subset selection — in a cluster of highly positively correlated actions, one may retain just one, without considerably reducing the performance.
1.1 Reward Approximation
@@
Now, suppose that the decision-maker has access to a certain subset of actions, $\mathcal{A}\subset\mathcal{A}_{\operatorname*{full}}$, rather than to the full set $\mathcal{A}_{\operatorname*{full}}$. For a bandit instance $\theta$, define the regret of being restricted to subset $\mathcal{A}$ as the gap between the best outcome achievable with the full action space and the best outcome within $\mathcal{A}$:
$\textnormal{Regret}(\theta):=\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta)-\max_{a^{\prime}\in\mathcal{A}}\;\mu_{a^{\prime}}(\theta),$} (1)
@@
Our objective then is to identify a small subset $\mathcal{A}$ that minimizes the expected regret $\mathbb{E}_{\theta\sim\mathcal{P}}\left[\operatorname*{Regret}\right]$, where the expectation is taken over all possible bandit instances.
This objective may be motivated in several ways: First, as discussed above, in problems such as inventory management it may be of natural interest to find a representative subset of actions, due to the intrinsic difficulties associated with maintaining the full set of actions.
Another possible reason would be to reduce the exploration phase when encountering a new bandit instance, especially when the time horizon $T$ is small. Indeed, the classic Bayesian bandit regret, (Agrawal & Goyal, 2012), generally scales as $\Omega(\sqrt{\left|\mathcal{A}_{\operatorname*{full}}\right|T})$, and may become prohibitively large for large $|\mathcal{A}_{\operatorname*{full}}|$. On the other hand, if a subset $\mathcal{A}$ is carefully chosen, the resulting Bayesian bandit regret can be significantly lower. To see this, for a given $\mathcal{A}\subset\mathcal{A}_{\operatorname*{full}}$ we can decompose the bandit regret as follows:
@@
$\textnormal{BayesianBanditRegret}:=\mathbb{E}\sum_{t=1}^{T}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta)-\mu_{A_{t}}(\theta)\right]$
$=\mathbb{E}\sum_{t=1}^{T}\left[\max_{a\in\mathcal{A}}\mu_{a}(\theta)\!-\!\mu_{A_{t}}+\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta)\!-\!\max_{a\in\mathcal{A}}\mu_{a}(\theta)\right]$
$\leq C\sqrt{|\mathcal{A}|\cdot T\log T}+T\cdot\mathbb{E}_{\theta}[\textnormal{Regret}].$ (2)
[…]
A further motivation is to reduce the computational burden that persists even in structured bandits; see Remark 4.8 for a discussion of linear and Gaussian process (GP) bandits.
1.2 Subset Selection Algorithm
We now introduce the subset selection algorithm itself, Algorithm 1. Conceptually, the algorithm is simple: Sample a bandit instance $\theta$ from a distribution $\mathcal{P}$ and solve it using an oracle (or practically, any of the standard bandit algorithms, such as UCB or TS), to find the optimal action $a^{*}=a^{*}(\theta)$. Add the action to the set or representation actions $\mathcal{A}$. Repeat the process $K$ times.
In the rest of the paper we show that the action subset generated by this simple algorithm will be diverse enough to appropriately cover $\mathcal{A}_{\operatorname*{full}}$ by $\mathcal{A}$. Our bounds will depend on the metric structure of the set $\mathcal{A}_{\operatorname*{full}}$ under the $L_{2}$ metric, and on the number of samples $K$. Note that Algorithm 1 does not require the knowledge of the metric, only the ability to sample bandit instances at random, and to solve them.
Algorithm 1 Epsilon Net Algorithm
2:Output: A subset of actions $\mathcal{A}$.
[…]
4:for $1,\dots,K$ do
5:  Sample a bandit instance $\theta$ from a distribution $\mathcal{P}$ (unknown to the learner).
6:  Find optimal action $a^{*}(\theta)\!:=\!\arg\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta)$.
7:  $\mathcal{A}\leftarrow\mathcal{A}\cup\{a^{*}\}$.
8:end for
Our main contributions are organized as follows:
  • Problem Formalization and An Algorithm Although we believe the subset selection problem is natural and of practical interest, to the best of our knowledge it has not been studied in the literature. In this paper we provide the first formalization of the problem, in the framework of the statistical similarities between actions. We also introduce a simple but effective subset selection algorithm, that does not require apriori knowledge of the correlations.
  • Regret Analysis. When the action sets exhibits sufficient correlation structure (ex. clustering or a low dimensional manifold structure), we provide upper and lower bounds on the expected error of the algorithm, i.e. the regret (1) (Section 4). The analysis is done under the standard assumption that the reward process $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ is Gaussian, (Srinivas et al., 2009; Garnett, 2023) and shows the basic relations between the induced geometry of the action set and the expected rewards. Our arguments can also be applied virtually with no changes to the sub-Gaussian processes, and this extension is carried out in Supplementary I. An extension of these arguments to other types of tail behavior may be implemented using standard techniques.
  • Generalization and Empirical Validation. We extend the analysis to settings where outcome functions lie in a reproducing kernel Hilbert space (RKHS). We empirically evaluate the subset selection algorithm on a number of problems (Section 5). Although there are no existing algorithms that could serve as direct performance comparison benchmarks, one could naively apply standard bandit algorithms, such as TS and UCB, via the notions of “super-arms”. As expected, since these methods are not sensitive to the underlying geometry of the subsets, we show that their performance is much weaker than that of our approach.
2 Related Work
For a general overview of multi-armed bandits (MAB), see (Lattimore & Szepesvári, 2020; Slivkins et al., 2019). Background on optimal action identification—which addresses the sample complexity of the inner optimization step in Algorithm 1—is provided in Supplementary A, along with regret bounds for structured (linear/GP) bandits.
While regret in general Bayesian bandits scales with $|\mathcal{A}_{\operatorname*{full}}|$, linear/GP bandits achieve $|\mathcal{A}_{\operatorname*{full}}|$-independent regret that depends on structural complexity (e.g., feature dimension or information gain). Their computational complexity, however, typically remains $|\mathcal{A}_{\operatorname*{full}}|$-dependent. Further discussion of computational aspects appears in Remark 4.8.
@@ 2 Related Work
Before discussing two related approaches that select a fixed-size subset $|\mathcal{A}|=K$–a choice often unclear in practice and requiring algorithm restarts–we highlight a key advantage of our method: it can adapt the subset cardinality on the fly.
@@
Top-K Action Identification aims to identify the $K$ actions with the highest expected outcomes using as few samples as possible (Kalyanakrishnan et al., 2012; Gabillon et al., 2012; Kaufmann et al., 2016; Chen et al., 2017). This line of work assumes that all actions are independent and have distinct expected values, making its methods inapplicable to our framework. If one were to apply these methods regardless, the most reasonable approach, in our view, would be to treat the family of bandit instances as a super-bandit, where each bandit instance corresponds to a round, and the expected payoff of an action in that round is given by $\mu_{a}$. In this setting, top-$K$ identification would refer to selecting $K$ actions with the highest expected payoffs $\mathbb{E}[\mu_{a}]$. In contrast, our framework considers that the expected outcome of each action, averaged over the distribution of bandits, may be the same—i.e., $\mathbb{E}[\mu_{a}]\!=\!c$ for all $a\!\in\!\mathcal{A}_{\operatorname*{full}}$, where $c$ is a constant. Further, even if $\mathbb{E}[\mu_{a}]$ varies across actions, ignoring correlations can be fatal in our framework:
Example 2.1.
@@
Combinatorial Bandits considers that the decision maker selects $K$ of base arms from $\mathcal{A}_{\operatorname*{full}}$ in each round, forming a super-arm $\mathcal{A}$, with $|\mathcal{A}|=K$. Popular methods include CUCB (Chen et al., 2016), CTS (Wang & Chen, 2018). We include this line of work as an empirical baseline, though it is not directly applicable to our framework for two reasons: (1) It assumes that the expected outcome of a super-arm depends only on the expected outcomes of its individual base arms, or imposes a stricter monotonicity condition. In our case, even though $\mathbb{E}[\mu_{a}]=c$ for all $a\in\mathcal{A}_{\operatorname*{full}}$, super-arm expected outcomes $\mathbb{E}[\max_{a\in\mathcal{A}}\mu_{a}]$ can differ significantly due to correlations among actions. (2) It assumes independence across base arms, whereas we explicitly model correlations. Ignoring these correlations misses the core challenge—an issue illustrated in Example 2.1.
3 Subset Selection Algorithm
In this Section, we introduce the formal problem setup and the technical tools used in the analysis.
The optimal action in a bandit instance $\theta$ is defined as
$a^{*}(\theta):=\arg\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta).$ (3)
Thus $a^{*}$ is a random variable taking values in the set $\mathcal{A}_{\operatorname*{full}}$. We denote by $q$ the induced probability distribution of $a^{*}$ on $\mathcal{A}_{\operatorname*{full}}$. Specifically, for any subset $r\subset\mathcal{A}_{\operatorname*{full}}$, we have
$q(r):=\mathbb{P}\left(a^{*}(\theta)\in r\right)=\mathbb{E}_{\theta\sim\mathcal{P}}\mathbbm{1}_{\{a^{*}(\theta)\in r\}}.$ (4)
In particular, note that the samples $a^{*}$ in lines 5&65\&6 of Algorithm 1 may be viewed as i.i.d samples from $q$. The distribution $q$ will be named as the importance measure $q$.
3.1 Gaussian Processes
As discussed earlier, in our analysis we consider the setting where the reward expectations $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ form a Gaussian process. An extension of our main bound to sub Gaussian processes is given in Supplementary I.
Let $\|\cdot\|_{2}$ denote the standard Euclidean norm on $\mathbb{R}^{n}$, and for a set $S\subset\mathbb{R}^{n}$ define the diameter of $S$ as $\mathrm{diam}(S):=\sup_{a,b\in S}\|a-b\|_{2}$.
To simplify the notation, we will consider the canonical representation of the Gaussian process: That is, we assume that $\mathcal{A}_{\operatorname*{full}}\subset\mathbb{R}^{n}$ for some dimension $n>0$, and that